OpenStudy (anonymous):

what does it mean in integration by parts when it says to move all the I terms to one side?

OpenStudy (anonymous):

Hmm, depends on the situation.

OpenStudy (anonymous):

Integration by parts is usually when you're trying to integrate the product of two functions.

OpenStudy (anonymous):

The formula is \[\int\limits u(x) \cdot dv = u(x) \cdot v(x) - \int\limits v(x) \cdot du\]

OpenStudy (anonymous):

well for the problem im on (sin4thets)(1/5)e^5theta-4/25(cos4theta)(e^5theta)-16/25 I+C(1)

OpenStudy (anonymous):

That looks like the result of an integration, what did you start with?

OpenStudy (anonymous):

I'm starting to think that this is one of those situations where one of the terms in your integration result is equal to the integral that you started with. Is that what you mean by I?

OpenStudy (anonymous):

yessss